OpenStudy (anonymous):

Thor invests $750 in an account, with interest compounded continuously. If his investment doubles in value after 9 years, how much money will he have after 12 years? [Use the formla A = Pert and round answer to the nearest cent.]

OpenStudy (anonymous):

2 = e^(9r) -> ln2 = 9r => r = ln2/9 Thus after 12 years: A = 750 * e^rt = $1889.88